OpenStudy (anonymous):

Some airlines have restrictions on the size of items of luggage that passengers are allowed to take with them. Suppose that one has a rule that the sum of the length, width and height of any piece of luggage must be less than or equal to 204 cm. A passenger wants to take a box of the maximum allowable volume. If the length and width are to be equal, what should the dimensions be? length = width =___________ height =________ In this case, what is the volume? volume =___________ (for each, include units) f the length is be twice the width, what should the dimensions be? length =

OpenStudy (anonymous):

make a cube

OpenStudy (anonymous):

Continuation width = height = In this case, what is the volume? volume = (for each, include units)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

after the box. whats next?

OpenStudy (anonymous):

@satellite73

OpenStudy (aum):

Volume V = L * W * H L + W + H = 204 L = W (given) V = W * W * H = W^2H 2W + H = 204 V = W^2(204 - 2W) Maximize V

OpenStudy (anonymous):

@aum i found that length = width =68cm height= 68 could help me get the volume?

OpenStudy (aum):

Yes. V = 68^3 cm^3

OpenStudy (anonymous):

can you help me with the second part too?

OpenStudy (aum):

\(\large V = 314432 ~~cm^3\)

OpenStudy (aum):

Follow the same procedure as the first.

OpenStudy (anonymous):

the second part is If the length is be twice the width, what should the dimensions be? length = width = height = In this case, what is the volume? volume = (for each, include units)

OpenStudy (aum):

Same procedure as the first: Volume V = L * W * H L + W + H = 204 L = 2*W (given) V = 2W * W * H = 2 * W^2 * H 3W + H = 204 V = 2W^2(204 - 3W) Maximize V

OpenStudy (anonymous):

@aum i found that height is 68cm

OpenStudy (aum):

Yes. What about L and W?

OpenStudy (anonymous):

im kinda lost there

OpenStudy (aum):

How did you find H without finding either L or W?

OpenStudy (anonymous):

what i did was i got this 2x+y = 204 then i solved for y and i got y= 204-2x (then i multiplied by x^2) and i got y = 204x^2 -2x^3 (then took the derivative) and i got 408x-6x^2 then solved for x i got x = 0 and 68, i discarded x = 0 and thats how i found height @aum

OpenStudy (anonymous):

that leaves me to think that there is 134cm to spare

OpenStudy (anonymous):

and that is all i got

OpenStudy (anonymous):

@aum you get it now?

OpenStudy (aum):

The above looks like answer for the first problem.

OpenStudy (aum):

If instead of L and W you want to use x and y that is fine. Then looks t my steps for the second part of the problem, replace L by y and W by x. There is still H for height and you can use z if you want.

OpenStudy (aum):

Let length = y width = x height = z Volume = xyz --- (1) x + y + z = 204 ---- (2) length is twice the width means y = 2x. Put this in (1) and (2) (1) becomes: Volume = xyz = x(2x)z = 2x^2z --- (3) (2) become: x + 2x + z = 204; 3x + z = 204; z = 204 - 3x ---- (4) Put (4) in (3): Volume V = 2x^2(204-3x) = 408x^2 - 6x^3 Maximize volume V dV/dx = 408*2*x - 18x^2 = 0. x(816 - 18x) = 0; x = 816/18 = 45.33 cm. Put x = 45.33 in (4) and find z. Then put x and z in (2) and find y.

OpenStudy (anonymous):

im lost afte the volume equation

OpenStudy (anonymous):

got it thanks again @aum